# Decentralization of Ethereum’s Builder Market

### v1

@@ Abstract.
Blockchains protect an ecosystem worth more than \$500bn with their strong security properties derived from the principle of decentralization. Is today’s blockchain really decentralized? In this paper, we empirically studied one of the least decentralized parts of Ethereum—the most used blockchain system in practice—and shed light on the decentralization issue from a new perspective.
To avoid centralization caused by Maximal Extractable Value (MEV), Ethereum adopts a novel mechanism that produces blocks through a builder market. After two years in operation, however, the builder market has evolved to a highly centralized one with three builders producing more than 90% of blocks. Why does the builder market centralize, given that it is permissionless and anyone can join? Moreover, what are the security implications of a centralized builder market to MEV-Boost auctions? Through a rigorous empirical study of the builder market’s core mechanism, MEV-Boost auctions, we answered these two questions using a large-scale auction dataset we curated since 2022.
Unlike previous works that focus on who wins the auctions, we focus on why they win, to shed light on the openness, competitiveness, and efficiency of MEV-Boost auctions. We show the access barriers around private order flows constitute a significant entry barrier to the builder market. A new builder needs to pay up to 1.4 ETH to access this private order flow, a significant amount considering their low initial profits. By computing the true value of bids, we investigated the inequality in block-building capabilities among builders and the competitiveness and efficiency of the MEV-Boost auctions. We observed that top, middle, and tail builders have significantly varying capabilities in extracting MEV, and the inequality worsens when MEV increases. 88% of the MEV-Boost auctions we studied were competitive, yet uncompetitive auctions still resulted in a total loss of 221.09 ETH for the proposers, which is 0.98% of their total gain. Meanwhile, only about 79% of the MEV-Boost auctions were efficient, and over half of the inefficient auctions were caused by block subsidization. Our findings also help identify directions for improving the decentralization of builder markets.
1. Introduction
@@
At the core of any decentralized blockchain, a key assumption underpinning their strong security properties (integrity, availability, censorship resistance, incentive compatibility, etc) is “decentralization”, that the system is run by a large number of independent participants (Nakamoto, 2008; Eyal and Sirer, 2018; Schwarz-Schilling et al., 2022; Wahrstätter et al., 2023a; Bahrani et al., 2024; Roughgarden, 2020). It is community consensus that one of the greatest threats to decentralization is Maximal Extractable Value (MEV) (Daian et al., 2020), profits that protocol participants, hereafter referred to as validators, can reap by manipulating the ordering of transactions. MEV can lead to centralization because it disproportionally benefits well-resourced validators compared to “regular” ones, as extracting MEV requires significant capital, intelligence, and computational resources, which regular validators may not have (Bahrani et al., 2024; Ethereum Foundation, 2024b; Pai and Resnick, 2023; Buterin, 2021b, a).
@@
In an effort to level the playground for validators, the Ethereum community proposed Proposer [^1] A proposer is a validator chosen as the leader in a round of consensus to add a new block to the blockchain.-Builder Separation (PBS) (Ethereum Foundation, 2024b). The idea is to outsource the task of building profitable blocks to specialized entities called block builders. In the current implementation called MEV-Boost (Flashbots, 2024), the proposer runs an auction among builders and adds the block with the highest bid to the blockchain. MEV-Boost is permissionless for builders so any builder can join the auction. Ideally, builders compete with each other in auctions, and the auctioneer—the proposer—collects a significant portion of MEV in auction revenue without doing much work, eliminating the competitive advantage of sophisticated validators over regular ones.
@@
Figure 1. Market share of builders. The builder market is arguably one of the least decentralized parts of Ethereum!
Does PBS achieve its goal? The answer is rather unclear. As shown in Fig. 1, the current market share is highly concentrated, with over 90% of the blocks produced through MEV-Boost being built by just three builders. However, the current concentration does not necessarily defeat the purpose of PBS for two reasons: first, builder markets may evolve and new builders may join and disrupt the concentration; second, centralization does not imply a lack of competition—it only takes a few non-colluding bidders to make the auction competitive and efficient.
Thus, while previous works (Flashbots, 2024; Wahrstätter, 2024; Labs, 2024; Heimbach et al., 2023; Wahrstätter et al., 2023b; Yang et al., 2022; Thiery, 2023) point out the centralization issue, they did not answer two key questions: First, how open is the builder market to new builders? Second, despite the centralization, are MEV auctions actually competitive and efficient? (We leave technical definitions to future sections, but intuitively, an auction is competitive if the winner outbids the second-best bidder; an auction is efficient if the bidder with the highest valuation wins.)
@@
In this paper, we answer these questions through a rigorous empirical study of historical MEV auctions. The key ingredient of our study is a large-scale auction dataset we curated since 2022: we collected 5.57 billion partial bids [^2] Partial and full bids are to be defined in Sec. 3.1 of 3,810,630 auctions from September 2022 to March 2024; in addition, in collaboration with ultra sound relay (Money, 2023), we obtained full bids for a total of 147,926 auctions from April 2023 to August 2023.
Auction data gives us direct visibility into the openness, competitiveness, and efficiency of MEV-Boost auctions. For example, historical bids allow us to “simulate” past auctions assuming a different set of transactions were available to builders, allowing us to discern the importance of different transactions. Another example is that we can compute the true value of each bid (i.e., the actual worth of the block to a builder, which is usually higher than what they bid) from transactions. The competitiveness and efficiency of an auction follow directly from a comparison between the true values and bid values. This is in contrast to, say, eBay auctions, where the true value of an item is subjective and cannot be retroactively computed from data.
Now we present an overview of methods and findings.
Table 1. Summary of findings
Finding 1: Private order flows contribute to 60% of the MEV in over 50% of the blocks in MEV-Boost.
Finding 2: To access private order flows, new builders need to pay up to 1.4 ETH to enter the market.
Finding 3: The inequality in block building is highest among tail builders, followed by middle and least among top builders, and it worsens when MEV increases.
Finding 4: 88.84% of the MEV-Boost auctions are competitive, with proposers’ losses from uncompetitive auctions amounting to 0.98% of total gains.
Finding 5: 79.74% of the MEV-Boost auctions are efficient, and 51.4% of the inefficient auctions are caused by block subsidization.
1.1. Openness of the Builder Market
First, we investigate the reasons behind centralization by quantifying the market entry barrier for a new builder. One such barrier could be the ability to build effective MEV extraction algorithms, but we investigated a perhaps more basic barrier: the ability to access profitable transactions. Recall that MEV extraction is about manipulating the ordering of transactions, so having access to “extractable” transactions is the basis of any extraction algorithm.
As we will present in Sec. 4, private order flows (Kilbourn, 2022)—the transactions that bypass the public mempool—contribute an increasing portion to builder income. This means that it would be difficult for a builder to win the auction without access to private order flows. Moreover, not all private order flows are equally profitable for builders, so the ability to access “good” private order flows forms a barrier to entering the builder market.
Using the auction dataset, we identified the providers of profitable transactions, quantified their importance to the builders, and investigated their accessibility to new builders.
Identifying pivotal private order flow providers.   In today’s MEV ecosystem (to be presented in Sec. 2.2), private order flows come from various private order flow providers. Different providers impose different access barriers to their private order flows. Some providers (such as MEV-Share (Flashbots, 2024) and MEV Blocker (MEV Blocker, 2023)) have relatively clear criteria a builder must satisfy to access their flows. Still, some providers (e.g., most searchers) are almost completely opaque to the public. To identify important providers, we define the pivotal level of a provider $P$ as the percentage of MEV-Boost auctions in which the winner would have lost had transactions from $P$ were not available. As we will show in Sec. 4.1, we identified five pivotal providers whose pivotal level exceeds 50% over a period longer than two weeks. That is, if a builder cannot access the flows from any of these providers, it will lose the majority of the auctions during the period of our study.
The good news is that two of these pivotal providers are MEV Blocker (MEV Blocker, 2023) and MEV-Share (Flashbots, 2024), which are open to builders who meet a relatively clear set of criteria. The bad news is that satisfying the requirements set by MEV Blocker and MEV-Share is non-trivial for new builders, which we discuss next.
Quantifying the market entrance barrier.   Typically, providers share private order flows to reputable builders with a certain market share (Sec. 4.2). However, this presents a chicken-and-egg problem for new builders: to gain market shares, they need private order flows to win the auction, but starting out new, they do not have market share and thus cannot access private order flows. To navigate this dilemma, new builders resort to subsidization (Flashbots, 2024): bidding higher than their actual profits in MEV-Boost auctions, to establish the required market shares. Therefore a new builder needs to have enough capital to keep subsidizing until reaching a certain market share. To understand the significance of this entry barrier, we compute the minimal cost of subsidization required to maintain a 1% market share, which is the requirement of MEV Blocker. Specifically, we first calculate the minimal subsidy a builder needs to win an MEV-Boost auction without private order flows; then, we sum up the minimum subsidies over 1% of total auctions per day as the daily cost.
Our analysis highlights a non-trivial cost for new builders to establish a reputation through subsidization—they need to invest up to 1.4 ETH to subsidize their blocks and secure a minimal 1% market share during the period of our study. The cost is comparable to or even higher than the profits of small builders (Flashbots, 2024).
Trust barriers result in access barriers.   An important takeaway is that siloes of private order flows stem from the lack of a fair exchange mechanism between providers and builders. In the current design, builders are at an advantage and malicious builders can harm providers by, e.g., sandwich attacks (Zhou et al., 2021) and imitation attacks (Qin et al., 2023). Therefore, providers only share their order flows with reputable builders, even though such reputation-based protection is pretty weak. As entry barriers arise from the trust crisis, reducing or eliminating these trust barriers is essential for making the builder market more decentralized. We discuss several ideas in Sec. 7.2.
1.2. Competitiveness and Efficiency
The first question focuses on the evolution of builder markets. In the second question, we focus on the status quo and study the implications of a centralized builder market as we have today.
MEV-Boost auction adopts an open-bid, ascending price auction akin to an English auction (Besanko et al., 2009), with a 12-second deadline. English auctions have appealing features: ideally, assuming bidders do not collude and can react to each other instantaneously, the bidders who can extract most MEV will win (efficiency), paying at least the MEV extracted by the second-highest bidder (competitiveness).
However, given the small number of builders in today’s builder market, it is unclear if MEV-Boost auctions indeed induce the desired dynamics. We investigate whether the apparent centralization of the builder market indeed leads to inadequate competition and a loss of efficiency.
True values in MEV-Boost auctions.   In MEV-Boost auctions, for a block $B$, a bid value (BV) is the amount the builder is willing to pay for $B$ to be added to the blockchain. The total revenue gained by the bidder from $B$ is the true value (TV). Rational bidders would bid lower than the true value to get a non-zero profit. The true value of a block $B$ can computed from transactions in $B$.
@@ 1.2. Competitiveness and Efficiency
Inequality in block building capabilities.   We first compare the true values across different builders for the same auction to understand the variations in builders’ abilities to extract values. A builder who consistently demonstrates a superior ability to extract value could dominate the market. We used the quartile coefficient of dispersion (QCD) (Wikipedia, 2022) to quantify the inequality of true values across builders. Our results in Sec. 5.1 reveal significant differences in the block-building capabilities of top, middle, and tail builders. The inequality worsens as the MEV of a slot increases, possibly because significant MEV opportunities are more likely to be shared with only a limited number of builders. Fortunately, the inequality among top builders is not significant, which supports the premise of competitiveness and efficiency.
More than 88% of auctions are competitive.   Using the true values, we propose two metrics, competitive index (CI) and efficient index (EI), to quantify whether an auction is competitive or efficient, respectively (see Sec. 5.2 for details.) Our results reveal that approximately 88.84% of the MEV-Boost auctions we analyzed were competitive. However, we also found that the auctions tend to be less competitive when the MEV of a slot increases. As discussed previously, the main concern is that a lack of competition can cause loss to proposers. We therefore quantify the loss for proposers. Our study shows that 16,498 of the 147,926 MEV-Boost auctions were uncompetitive, where proposers incurred total losses amounting to 221.09 ETH, which is 0.98% of the total gains of the proposers.
Only 79% of auctions are efficient.   An interesting finding is that more than 20% of MEV-Boost auctions are inefficient and the auctions tend to be inefficient when the MEV in an MEV-Boost auction is low. Further analysis reveals that over half of the inefficient auctions are caused by block subsidization. The winners of these inefficient auctions overbid to win, although they do not have the highest true value. Moreover, block subsidization is more common when the extractable value is small, because the cost will be relatively low. Although this may not benefit the overall welfare, it is advantageous for proposers as they earn more than expected.
Suggestions for future developments.   Since many identified issues stem from trust concerns, future developments should aim to design mechanisms that reduce or eliminate the reliance on trust. Possible directions include community supervision (Flashbots, 2023a) and the use of Trusted Execution Environments (TEEs) (Costan and Devadas, 2016) to ensure the integrity and confidentiality of private order flows (Flashbots, 2024c). Another idea is to implement a reliable leakage detection mechanism for private order flows, which we briefly discuss in Sec. 7.2. Such mechanisms would encourage private order flow providers to distribute their flows across multiple builders, thus increasing the equality of the builders’ ability to extract value.
@@
To the best of our knowledge, we are the first to use the rich auction data to analyze the decentralization of Ethereum’s builder market, while most previous works are limited to on-chain data. Our dataset of the true values of 147,926 MEV-Boost auctions from April to August 2023 may be of independent interest. Our dataset of 5.57 billion partial bids since 2022 is the largest-scale known to us.
We characterized the efficacy of PBS in terms of being open to new builders and inducing competitive and efficient auctions. Our key findings are highlighted in Tab. 1, which we summarize below.
  • Sec. 4: Private order flows contribute to more than 60% of the MEV in over 50% of blocks every day since July 2023. We identified five pivotal private order flow providers that had a significant and sustained impact on the winners of more than half of the MEV-Boost auctions. Two of them are MEV Blocker and MEV-Share.
  • Sec. 4.2: Private order flow providers impose market share requirements on builders, which costs new builders up to 1.4 ETH in subsidy to enter the market during the period of our study.
  • Sec. 5.1: The QCD shows a significant inequality in builders’ abilities to extract value across top, middle, and tail builders, but the inequality among the top builders is not significant.
  • Sec. 5.2: We find that 88.84% of the 147,926 MEV-Boost auctions are competitive. However, proposers incur losses of 221.09 ETH from the uncompetitive auctions, representing 0.98% of their total gain. These losses can still threaten the stable separation between proposers and builders.
  • Sec. 5.2: 79.74% of the MEV-Boost auctions in our study are efficient. A significant factor contributing to inefficiency is block subsidization, where winners overbid to secure the auctions, possibly driven by the need to meet market share requirements.
@@
Besides contributing to the literature on MEV mitigation, our work also contributes to the Economics literature on the analysis of MEV-Boost auctions. Currently, there is only limited auction theoretic analysis that considers the complexity of real-world MEV auctions (Gupta et al., 2023; Pai and Resnick, 2023; Wu et al., 2024). The findings in this paper can serve as empirical evidence for future theoretic analysis of MEV-Boost auctions.
Figure 2. Illustration of the MEV supply chain in the PBS scheme.
2. Background
2.1. MEV-Boost Auctions

### v3

@@ Abstract
Blockchains protect an ecosystem worth more than \$500bn with strong security properties derived from the principle of decentralization. Is today’s blockchain decentralized? In this paper, we empirically studied one of the least decentralized parts of Ethereum, its builder market.
The builder market was introduced to fairly distribute Maximal Extractable Values (MEV) among validators and avoid validator centralization. As of the time of writing, three builders produced the vast majority (more than 80%) of blocks in Ethereum, creating a concerning centralization factor. However, the community believes that such centralization “is okay,” arguing that builder centralization will not lead to validator centralization. In this empirical study, we interrogate the causes and implications of builder centralization and challenge this belief that it is acceptable.
Our main finding is that builder centralization has led to a significant loss by validators and, if left uncontrolled, could lead to validator centralization. Moreover, MEV mitigation solutions slated for adoption are affected too because they rely on the builder market as an “MEV oracle,” which is made inaccurate by centralization.
Our investigation revealed two reasons behind builder centralization. We propose a structural change to the existing MEV supply chain and a solution based on the new supply chain structure. However, future work is required to analyze if the new supply chain structure is sustainable in the long term, which we leave open.
1 Introduction
@@
At the core of any decentralized blockchain, a key assumption underpinning their strong security properties (integrity, availability, censorship resistance, incentive compatibility, etc.) is “decentralization”, that the system is run by a large number of independent participants [63, 28, 75, 83, 4, 74]. It is community consensus that one of the greatest threats to decentralization is Maximal Extractable Value (MEV) [16], profits that protocol participants, known as validators, can reap by manipulating the ordering of transactions. MEV can lead to centralization because it disproportionally benefits well-resourced validators compared to “regular” ones, as extracting MEV requires significant capital, intelligence, and computational resources, which regular validators may not have [4, 23, 67, 12, 11].
@@
To distribute MEV fairly among validators and avoid validator centralization, Ethereum adopted a solution called Proposer-Builder Separation (PBS) [23]. A proposer is the validator chosen to add a new block to the blockchain. In the current implementation called MEV-Boost [35], the proposer runs an auction among builders (anyone can be a builder) to solicit the best block. Ideally, builders compete with each other, and the proposer collects a significant portion of MEV in auction revenue without doing much work, eliminating the advantage of sophisticated validators over regular ones.
@@
Figure 1: Market share of builders. The builder market is one of the least decentralized parts of Ethereum!
Is the builder market working as intended? The answer is rather unclear because the builder market is highly centralized. As shown in Fig. 1, three builders won over 90% of auctions, raising questions about the competitiveness of the market. However, the community believes that “it is OK to centralize block building” [23], arguing that block centralization will not lead to validator centralization. We challenge this belief through an empirical study of the past MEV-Boost auctions. As will be shown in detail, builder centralization has caused a significant loss to the proposers and, if left uncontrolled, could lead to undesired consequences including validator centralization.
1.1 Implications of Builder Centralization
Proposer loss and its consequences.   Intuitively, a proposer’s loss is the difference between the MEV available in a block and the auction revenue. Builder centralization can cause proposer loss in two ways. First, auctions may not be competitive enough due to, e.g., collusion or other limitations of the auction mechanism. Second, even if builders compete with each other, if a single builder has a unique advantage, it will not be incentivized to bid on all of the MEV it can extract. Instead, it only needs to outbid the second-best bidder slightly to win. We say that the proposer incurs a loss from uncompetitiveness in the first case and a loss from inequality in the second case.
If proposers suffer significant losses, they would be incentivized to extract MEV by themselves (instead of using the builder market). This is undesirable because it re-introduces the dynamics that PBS wants to avoid: sophisticated validators would have an advantage over small ones in extracting MEV. In other words, significant proposer losses would make PBS “unstable”. Moreover, several MEV mitigation solutions (e.g., MEV-burn [19]) rely on the MEV-Boost auctions as an “oracle” to know how much MEV has been extracted from the system (e.g., MEV-burn proposed to remove that amount from circulation, effectively redistributing them to all holders of ETH through deflation.) The proposer’s loss translates to inaccuracy in the MEV oracle.
Next, we introduce the auction mechanism and present our method to quantify the proposer losses using auction data.
Modeling MEV-Boost auctions.   MEV-Boost auction adopts an open-bid, ascending price auction akin to an English auction [6], with a 12-second deadline. For each auction instance, builders submit bids in the form of $(B,BV)$ specifying a block $B$ and a bid value (BV) the builder is willing to pay for $B$ to be added to the blockchain. The revenue a bidder gains from $B$ is called the true value ($TV(B)$), which can be estimated from transactions in $B$. At the end of an auction, the block $B$ with the highest bid is added to the blockchain, the builder pays $BV$ to the proposer and earns $TV$ when $B$ is executed (i.e., the builder’s profit is $TV-BV$).
Recall that the proposer’s loss is the difference between the auction revenue and the MEV available in a block. The latter can be represented by the builders’ highest true values. With that, the two types of losses can be computed from auction data, as illustrated in Fig. 2. Specifically, suppose the true values of the bids from a given auction instance are $TV_{1}>TV_{2}\geq\cdots\geq TV_{N}$. The loss from inequality is $TV_{2}-TV_{1}$ because a competitive auction can, at most, generate revenue of $TV_{2}$. The loss from uncompetitiveness is $\max(TV_{2}-BV,0)$ (if this value is negative, the auction is considered competitive, and there is no loss from uncompetitiveness.)
Figure 2: Definition of proposer losses.
@@ 1.1 Implications of Builder Centralization
Auction data.   A practical challenge in quantifying the proposer loss in past MEV-Boost auctions is the lack of data. Only the winning bids are recorded in the blockchain, but to compute the losses, we also need losing bids. The key ingredient that enabled our study is a large-scale auction dataset we curated since 2022: we collected 8.79 billion partial bids [^1] Partial and full bids are to be defined in Sec. 3.1 of auctions from September 2022 to July 2024; in addition, in collaboration with ultra sound relay [62], we obtained full bids for a total of 301,479 auctions from April 2023 to December 2023. From the auction dataset, we can compute the true value of all bids.
Results.   91.8% of the MEV-Boost auctions we analyzed were competitive. However, auctions tend to be less competitive when the MEV of a slot increases. The loss from uncompetitiveness turns out rather insignificant, ranging from 0.5% to 1.7% during the periods of our study.
@@
To understand the variations in builders’ abilities to extract values, we statistically compared the true values across different builders in the same auction. We used several metrics, including the quartile coefficient of dispersion (QCD) [88], to quantify the inequality of true values across builders. We found significant differences in the block-building capacities of top, middle, and tail builders. The inequality worsens as the MEV of a slot increases, likely because significant MEV opportunities are shared with only a limited number of builders.
Inequality of block-building capacities has led to a proposer loss ranging from 5.6% to 11.5%. I.e., had there been no inequality, the proposer would have earned roughly 10% more. This indicates that inequality is the primary driver of the proposer’s losses in the MEV-Boost auctions.
1.2 What caused Inequality in Block-building Capacities
Given that proposers suffer significant losses due to the builders’ block-building capacity inequality, we now investigate the underlying causes of this inequality.
In the current MEV ecosystem (to be presented in Sec. 2.2), builders receive order flows from order flow providers and compute the optimal ordering of received order flows (see Appx. E for a reference builder algorithm). Thus, a builder’s block-building capacity is determined by the “quality” of the order flows it receives. At the same time, not all order flows are accessible to all builders. The difference in builder’s ability to access important private order flows drives the inequality in block-building capacity. This also explains builder centralization: only a small set of builders can access these profitable order flows.
We proposed a metric to identify important order flows, and we investigated their accessibility by different builders and the underlying reasons.
Identifying pivotal private order flow providers.   To formally capture the intuition of “important providers”, we defined pivotal level to quantify the sustained impact of a given provider. A provider $P$ is pivotal in an auction $A$ if the winner of $A$ would have lost, had transactions from $P$ were not available. In other words, pivotal providers are necessary for the winner to win. The pivotal level of a provider $P$ is the percentage of MEV-Boost auctions in which $P$ is pivotal.
Using the auction dataset, we identified five pivotal providers whose pivotal level exceeds 50% over a period longer than two weeks. That is, if a builder cannot access the flows from any of these providers, it will lose the majority of the auctions during our study period.
These order flows are not equally inaccessible to all builders for two reasons we now introduce.
Inequality due to trust barriers.   Order flow providers do not openly share private order flows with any builder because a malicious builder can harm the provider by, e.g., sandwich attacks [94] and imitation attacks [71]. Providers set up different access barriers for self-protection.
Most MEV searchers (a specific type of order flow providers) resort to “obscurity for security” in that they are not publicly reachable; thus, we could not learn about their access barriers and how to overcome them. On the other hand, MEV channels (another type of provider) typically publish a clear access policy, which allowed us to gain some insight into the nature of their access barrier.
For example, MEV-Blocker only shares order flows with builders with a market share of at least 1%, using market share as a proxy for reputation and trustworthiness. This requirement introduces a chicken-and-egg problem for new builders: to gain market share, they need access to private order flows to win auctions, but as newcomers, they lack the required market share to access these flows. As a result, new builders often resort to subsidization: bidding higher than their actual profits in MEV-Boost auctions to establish the required market shares. Therefore, a new builder needs to keep subsidizing until reaching a certain market share to overcome the access barriers set by MEV channels. The cost to do so continues to increase over time and can reach up to 7.4 ETH, according to our quantification.
In summary, with a moderate cost, builders can overcome the access barrier of one type of order flow provider (MEV channels). However, how to access order flows from other providers, particularly important searchers, remains unclear.
Inequality due to integration.   A more worrisome finding is that some providers send order flows almost exclusively to their preferred builders, causing an inequality in block-building capacities.
To identify such provider-builder pairs, we introduced builder-specific pivotal level of a provider $P$ for a builder $B$ as the pivotal level of $P$ in auctions won by $B$. With this metric, we rediscovered known exclusive relationships (called “integration” [40]), but also found unreported integrated providers for top builders (Titan and Beaver respectively). This finding is interesting because Titan has been the only top builder without a well-known integrated provider. Our findings show that all three top builders, accounting for 90% of market share, have integrated providers.
Integration poses a market entry barrier as new builders likely lack integrated providers. Moreover, it offers a unique competitive edge and will likely cause dominance. Finally, there is a strong incentive to integrate: compared to sending valuable order flows to multiple builders, sending them to a single builder enables this builder to win the auction with higher profit; the provider and the builder can share the profit, which means integration improves their total utility.
Integration is detrimental to proposers, as it will reduce their profits. The strong incentive for integration makes it increasingly difficult to address builder centralization.
1.3 Mitigation Ideas
To summarize, we found that trust barriers and incentives for integration contribute to builder centralization, leading proposers to incur significant losses. The consequences include the instability of PBS and the inaccuracy of MEV oracles.
We propose a structural change to the existing MEV supply chain (Sec. 2.2) and a solution based on the new supply chain structure. However, future work is required to analyze if the new supply chain structure is sustainable in the long term, which we leave open.
The new supply chain has only one builder to which the order flow providers send flows and from which the proposers accept blocks. In this setting, a recent work [57] suggests a truthful mechanism that incentivizes order flow providers to report the MEV in their order flows truthfully.
Our solution involves having the builder run the truthful mechanism in [57] in a Trusted Execution Environment (TEE) [15], hardware-based enclaves that guarantee the confidentiality and integrity of programs running inside. Order flow providers verify that the correct mechanism is running in the TEE using remote attestation, then send order flows to the TEE along with a bid indicating the MEV in their order flow. Then, the TEE computes a fair refund to order flow providers. The refund mechanism is designed so that bidding truthfully is a dominant strategy [57]. (While oversimplifying, one can make a loose analogy to sealed-bid second-price auctions among order flow providers.) Then, the TEE can compute the total MEV in a block from the providers’ bids, solving the problem of inaccurate MEV oracles.
This also addresses the instability of PBS because the builder program in the TEE can be programmed to pay the proposer fairly, leaving just enough for its operation. Since this builder is fair to order flow providers, it is conceivable that it will be preferred by order flow providers over other builders. However, as mentioned above, it is unclear if having a single builder is an equilibrium. Also, the proposers may be incentivized to become order flow providers, which may re-introduce the concern similar to PBS instability. We elaborate on the idea in Sec. 7.1.
@@ 1.3 Mitigation Ideas
To the best of our knowledge, we are the first to use the rich auction data to analyze the decentralization of Ethereum’s builder market, while most previous works are limited to on-chain data. We summarize our contribution below.
  • We found that builder centralization has led to significant proposer losses, which could lead to undesirable consequences, including validator centralization and inaccurate MEV oracles. We challenge the community’s belief that “builder centralization is okay".
  • We attribute proposer losses to two factors: auctions are not competitive, builders’ block-building capacities are unequal, and inequality is the primary cause.
  • We identify two reasons for inequality in block-building capacity: providers’ trust concerns and the incentive for integration.
  • Our dataset of the true values of 301,479 MEV-Boost auctions from April to December 2023 may be of independent interest. Our dataset of 8.79 billion partial bids since 2022 is the largest scale known to us.
@@
  • Besides contributing to the literature on MEV mitigation, our work also contributes to the Economics literature on the analysis of MEV-Boost auctions. Currently, there is only limited auction theoretic analysis that considers the complexity of real-world MEV auctions [40, 67, 90]. The findings in this paper can serve as empirical evidence for future theoretical interrogation.
Figure 3: Illustration of the MEV supply chain in the PBS scheme.
2 Background
2.1 MEV-Boost Auctions